$SPAD/src/input seg.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool seg.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
s := 3..10
--R
--R
--R   (1)  3..10
--R                                               Type: Segment(PositiveInteger)
--E 1
--S 2 of 10
lo s
--R
--R
--R   (2)  3
--R                                                        Type: PositiveInteger
--E 2
--S 3 of 10
hi s
--R
--R
--R   (3)  10
--R                                                        Type: PositiveInteger
--E 3
--S 4 of 10
t := 10..3 by -2
--R
--R
--R   (4)  10..3 by - 2
--R                                               Type: Segment(PositiveInteger)
--E 4
--S 5 of 10
incr s
--R

--R
--R   (5)  1
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 10
incr t
--R
--R
--R   (6)  - 2
--R                                                                Type: Integer
--E 6
--S 7 of 10
l := [1..3, 5, 9, 15..11 by -1]
--R
--R
--R   (7)  [1..3,5..5,9..9,15..11 by - 1]
--R                                         Type: List(Segment(PositiveInteger))
--E 7
--S 8 of 10
expand s
--R
--R
--R   (8)  [3,4,5,6,7,8,9,10]
--R                                                          Type: List(Integer)
--E 8
--S 9 of 10
expand t
--R
--R
--R   (9)  [10,8,6,4]
--R                                                          Type: List(Integer)
--E 9
--S 10 of 10
expand l
--R
--R
--R   (10)  [1,2,3,5,9,15,14,13,12,11]
--R                                                          Type: List(Integer)
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

5